A chemist wants to mix a 22% acid solution with a 36% acid solution to get 28 L of a 26% acid
solution. How many liters of the 22% solution and how many liters of the 36% solution should be
mixed?

Answer by richwmiller(17219)

a+b=28,
0.22*a+0.36*b=0.26*28
a=28-b
0.22*(28-b)+0.36*b=7.28
6.16-0.22b+0.36*b=7.28
0.14*b=1.12
b=8
a=28-b
a=20 liter at 22% b=8 liter at 36%
check
0.22*20+0.36*8=0.26*28
4.4+2.88=7.28
7.28=7.28
ok